OpenStudy (anonymous):

Integrate csc^4 x dx

OpenStudy (anonymous):

This might help get you started on the right path the integral of csc^2x= - cot

OpenStudy (anonymous):

and csc^2x = 1 + cot^2x

rishavraj (rishavraj):

do u mean \[\csc ^{4} x\]

rishavraj (rishavraj):

if its so then its \[\int\limits \csc ^{2} x \csc ^{2} x = \int\limits (1 + \cot ^{2}x) \csc ^{2}x\]

OpenStudy (anonymous):

\[-\frac{1}{3} \cot (x) \left(\csc ^2(x)+2\right) \]